# Understanding Long Videos with Multimodal Language Models

### v2

@@ 5.2 Robotics Domain Action Recognition
In this section, we investigate generalization capabilities of our proposed MVU by evaluating across various datasets from robotics domain Open X-Embodiment (Open-X-Embodiment-Collaboration et al., 2023), following a QnA style formulation of the dataset. See Appendix D for more details on this dataset. We highlight the visual differences of this data in Figure 4.
Our results on these robotics datasets are presented in Table 4 and indicate clear performance improvements for our proposed MVU over the baseline from Zhang et al. (2023a). The purpose of this experiment is to evaluate the generality of our approach to video domains different from everyday natural videos. We take these promising results to indicate the strong generality of our framework.
Furthermore, we note how our modality constrained variants do not perform significantly better than random on these robotics domain tasks (details in Appendix E). We attribute this to the significant domain shift in terms of the world of operation in this domain (i.e. robotics tasks tend to involve controlled environments very different to what humans face on an everyday basis).
Table 4: OpenX Detailed Results: We report accuracy (%) for the VideoQA formulation of Open X-Embodiment benchmark (Open-X-Embodiment-Collaboration et al., 2023). Our proposed MVU achieves clear improvements over random selection and LLoVi baseline (Zhang et al., 2023a). In the table header, Obs. (observation), size, CC (class count) stand for the camera used, number of videos, and number of unique language instructions per dataset, respectively. In the observation column, T stands for third-person view (stationary camera that does not move with robot), while F denotes first-person view where camera is mounted on moving robot. Note that total is average weighted by dataset size.
@@ 5.3 Ablations
In this section, we systematically dissect our overall MVU framework to establish the usefulness of each of its individual component. We first ablate our three different information modalities and report these results in Table 5. Our evaluations indicate clear performance boosts from each of our object-centric information modalities.
@@
We next ablate our overall framework against the existing work Zhang et al. (2023a) by replacing our MVU object-centric information pipeline with the frame description approach in Zhang et al. (2023a). We construct a setup identical to our framework except for the inputs to our final stage VLM replaced with frame level descriptions. These results reported in Table 7 indicate the clear significance and improvement of our proposed object-centric information pipeline over simple frame descriptions. The 8 frame variant is the same speed comparison as MVU uses captioner only on 8 frames. Our MVU outperforms both that and the slower 16 frame baseline. We also note the performance drop in the baseline when increasing the number of frames from 16 to 180. While consistent with observations in prior works for long-video tasks (Mangalam et al., 2023), we attribute this drop to decreased signal to noise ratio with introducing of additional frame descriptions. This further highlights the importance of selecting useful information from video frames and we reiterate how the object-centric information in our MVU framework serves this purpose.
@@ 6 Conclusion
Our method utilizes multiple pretrained models, all of which are open-source with model weights freely available. We use the versions of these models hosted on HuggingFace https://huggingface.co for all our experiments. We discuss all steps in our proposed algorithms in detail while also releasing relevant code. All evaluations we perform are on public datasets accessible by all (some behind evaluation servers to prevent test set contamination). Our code will also be released publicly.
@@ Appendix B Details on Pretrained Models and Datasets
Models: Our framework utilizes three distinct off-the-shelf models for its various operations, namely a) an LLM / VLM for likelihood selection (including visual frame selection and textual answer selection), b) a generative VLM for extracting object list from a frame, and c) an open-vocabulary detector for object localization. We use LLaVA-v1.5-13B (Liu et al., 2023a) for likelihood selection, frame selection, and frame object list generation. For object localization, we use OWL-ViT-B/32 (Minderer et al., 2022). Unless explicitly specified, we use the above setup in all our experiments. Variants of our framework uses LLMs Llama-2-7b-Chat, Gemma-7b-IT, and Mistral-7B-Instruct (default) for likelihood selection. Apart from these off-the-shelf models, our framework involves zero additional training. We also reiterate that no components of our framework undergo any form of video-level training.
@@ Appendix C Details on Baselines
In Section 5.1, we evaluate the performance on long-video understanding tasks using work in Zhang et al. (2023a) and Wang et al. (2023) as two baselines for comparison. However, both these methods utilize closed-source, proprietary LLMs (i.e. GPT-4) with parameter counts on the order of trillions (over 100X our model size) deeming their direct comparisons unfair. In the case of Zhang et al. (2023a), we replicate their method (using their open-source repository and pre-trained models following Kahatapitiya et al. (2024)) utilizing an open-source LLM of comparable parameter count as our framework. For Wang et al. (2023), the authors directly report results for a variant with a similar parameter count as ours. We utilize these evaluations as our point of comparison.
@@ F.1 Implementation Details
Appendix G Open-Ended Video Question Answering
Table A.5: Open-Ended Video QnA Evaluation: We present results on the ActivityNet dataset Yu et al. (2019) that demonstrate strong performance of our proposed MVU framework. Accuracy (%) is reported. VT stands for video level training. We highlight how our MVU framework utilizes no video level training for any of its components and surpassed multiple approaches that rely on video-language training.
@@ Appendix G Open-Ended Video Question Answering
Appendix H Additional Ablations
Therein we investigate a new motion focused dataset, Something-Something-v2 Goyal et al. (2017) (SSv2), only for this single ablation. The SSv2 dataset focuses on motion-based category discrimination, providing an ideal evaluation to measure the usefulness of our object motion trajectory modality. We benchmark on a subset of this dataset following Lin et al. (2023b) and report these results in Table A.6. Our results while exceeding their performance also indicate the clear performance gains obtained when injecting the object motion trajectory modality into our MVU framework.
@@ Appendix H Additional Ablations
Appendix I Tokenization in LLMs
Most modern LLMs utilize Byte-Pair Encoding (BPE) tokenization Sennrich et al. (2015) to convert natural language into discrete tokens that are mapped to vocabulary embeddings. This process is learned from language itself and the resulting tokenization may sometimes break complete words into pieces (e.g. example $\rightarrow$ ex-am-ple). Given our utilization of logits extracted from an LLM forward pass, we note that each logit corresponds to a single token, which may at times be the embedding of some meaningless word piece. However, our calculation of a joint likelihood across a sequence of tokens ensures a meaningful process, which is validated by our strong results.

### v3

@@ 5.2 Robotics Domain Action Recognition
We investigate generalization capabilities of our proposed MVU by evaluating across datasets from robotics domain Open X-Embodiment (Open-X-Embodiment-Collaboration et al., 2023), following a QnA style formulation of the dataset (details in Appendix D). We highlight visual differences of this data in Figure 4. We present evaluations in Table 4, which indicate clear performance improvements for MVU over the baseline from Zhang et al. (2023a). The purpose of this experiment is to evaluate the generality of our approach to video domains different from everyday natural videos. We take these promising results to indicate the strong generality of our framework. Furthermore, we note how our modality constrained variants do not perform significantly better than random on these robotics domain tasks (details in Appendix E). We attribute this to the significant domain shift in terms of the world of operation in this domain (i.e. robotics tasks tend to involve controlled environments very different to what humans face on an everyday basis).
Table 4: OpenX Detailed Results: We report accuracy (%) for the VideoQA formulation of Open X-Embodiment benchmark. MVU achieves clear improvements over random selection and LLoVi baseline (Zhang et al., 2023a). In table header, Obs. (observation), size, CC (class count) stand for camera used, number of videos, and number of unique language instructions per dataset, respectively. In observation column, T stands for third-person view (stationary camera that does not move with robot), while F denotes first-person view where camera is mounted on moving robot. Note that total is average weighted by dataset size.
@@ 5.3 Ablations
In this section, we systematically dissect our overall MVU framework to establish the usefulness of each of its individual component (see Appendix I for more ablations). We first ablate our three different information modalities and report these results in Table 5. Our evaluations indicate clear performance boosts from each of our object-centric information modalities.
@@
We next ablate our overall framework against the existing work, Zhang et al. (2023a), by replacing our MVU object-centric information pipeline with the frame description approach in Zhang et al. (2023a). We construct a setup identical to our framework except for the inputs to our final stage VLM replaced with frame level descriptions. These results reported in Table 7 indicate the clear significance and improvement of our proposed object-centric information pipeline over simple frame descriptions. The 8 frame variant is the same speed comparison as MVU uses captioner only on 8 frames. Our MVU outperforms both that and the slower 16 frame baseline. We also note the performance drop in the baseline when increasing the number of frames from 16 to 180. While consistent with observations in prior works for long-video tasks (Mangalam et al., 2023), we attribute this drop to decreased signal to noise ratio with introducing of additional frame descriptions. This further highlights the importance of selecting useful information from video frames and we reiterate how the object-centric information in our MVU framework serves this purpose.
@@ 6 Conclusion
Our method utilizes multiple pretrained models, all of which are open-source with model weights freely available. We use the versions of these models hosted on HuggingFace https://huggingface.co for all our experiments. For LVNet baseline, we utilize code from https://github.com/jongwoopark7978/LVNet. We discuss all steps in our proposed algorithms in detail while also releasing relevant code. All evaluations we perform are on public datasets accessible by all (some behind evaluation servers to prevent test set contamination). Our code will also be released publicly.
Contributions
KR led the project by building the preliminary ideas followed by performing most experiments and evaluations. XL contributed to ideas on off-the-shelf model usage, setup all robotics evaluations, debugged many issues, and discussed all aspects of the project. KK contributed to ideas on VLM prompting and templating of textual data, setup several baseline replications, reviewed code, and discussed all aspects of the project. MR organized the project, set the research direction, and discussed all aspects of the project idea, scope, and implementation.
Acknowledgments
We would like to thank all members of the Robot Learning Lab at Stony Brook University for support, feedback and guidance. In particular, we would like to thank Jongwoo Park for technical feedback and support; Si Chen for logistical help and encouragement throughout the project.
@@ Appendix B Details on Pretrained Models and Datasets
Models: Our framework utilizes three distinct off-the-shelf models for its various operations, namely a) an LLM / VLM for likelihood selection, b) a generative VLM for extracting object list from a frame, and c) an open-vocabulary detector for object localization. We use LLaVA-v1.5-13B (Liu et al., 2023a) for likelihood selection and frame object list generation. For object localization, we use OWL-ViT-B/32 (Minderer et al., 2022). Unless explicitly specified, we use the above setup in all our experiments. Variants of our framework uses LLMs Llama-2-7b-Chat, Gemma-7b-IT, and Mistral-7B-Instruct (default) for likelihood selection. Apart from these off-the-shelf models, our framework involves zero additional training. We also reiterate that no components of our framework undergo any form of video-level training.
@@ Appendix C Details on Baselines
In Section 5.1, we evaluate the performance on long-video understanding tasks using work in Zhang et al. (2023a) and Wang et al. (2023a) as two baselines for comparison. However, both these methods utilize closed-source, proprietary LLMs (i.e. GPT-4) with parameter counts on the order of trillions (over 100X our model size) deeming their direct comparisons unfair. In the case of Zhang et al. (2023a), we replicate their method (using their open-source repository and pre-trained models following Kahatapitiya et al. (2024)) utilizing an open-source LLM of comparable parameter count as our framework. For Wang et al. (2023a), the authors directly report results for a variant with a similar parameter count as ours. We utilize these evaluations as our point of comparison.
We highlight that re-implementations of these baselines utilize common LLMs / VLMs as our MVU framework followed by identical evaluation protocols to ensure fair comparison.
@@ F.1 Implementation Details
F.2 Distinction from Exact Match
As described in the previous section, likelihood selection uses a likelihood measure which is the likelihood (probability) of the model generating the given sentence (as opposed to being an exact match). This likelihood measure is also used as the training loss when training LLMs. Given how LLMs trained with this loss (i.e. all decoder based LLMs such as LLaMA, Gemini, GPT) are highly effective at handling semantic meaning, it follows that this loss can capture semantic meaning. This likelihood measure is calculated within the LLM latent space. This is equivalent to the probability (or likelihood) of that answer being generated by the LLM conditioned on the input question. We derive this in detail in Appendix F. Relating to the same example, this means that likelihood is an estimate of how likely the model would predict ‘C is washing plates’ as opposed to making that exact match. This means predictions closer to the target such as ‘C is cleaning dishes’ would also gain high likelihood values.
In fact, we validate this second point through a toy example. We provide an LLM with the question ”X is cleaning dishes in the kitchen. What is X doing? a) washing plates, b) cleaning laundry, c) painting dishes. The correct choice is:” and calculate the likelihood for each of the 3 responses. The calculated likelihoods are 0.996, 0.006, 0.007 for a, b, c respectively (highest is selected), despite response (a) having no common words with the original statement unlike (b) and (c). This illustrates the ability of likelihood selection to capture semantic meanings.
F.3 Detailed Prompting Example
We also note that while different choices are repeated along the batch, our likelihood implementation actually follows prior approaches where all answer candidates are fed together to the language model in addition to organizing the Q-A pairs in a batch dimension. Taking one simplified toy example, given a question “Where is the dog?” and answers “mat, table, bench”, we use three queries along batch dimension as:
  • Where is the dog? Select the correct response from: a) mat, b) table, c) bench. The correct response is a) mat.
  • Where is the dog? Select the correct response from: a) mat, b) table, c) bench. The correct response is b) table.
  • Where is the dog? Select the correct response from: a) mat, b) table, c) bench. The correct response is c) bench.
In fact, applying likelihood selection without such prompting leads to significantly low performance for some datasets. We show this in Table 7 which we repeat here as Table A.5.
Table A.5: Ablating Answer Candidates in Prompt: We illustrate the importance of appropriate prompting when combining with likelihood selection, specifically for long video QnA tasks. Top-1 accuracy (%) is reported on EgoSchema subset (ES-S) and NextQA test set (NQA-T).
Dataset ES-S NQA-T
No answer candidates in prompt 58.2 35.8
With answer candidates in prompt 60.3 55.4
Table A.6: Open-Ended Video QnA Evaluation: We present results on the ActivityNet dataset (Yu et al., 2019) that demonstrate strong performance of our proposed MVU framework. Accuracy (%) is reported. VT stands for video level training. We highlight how our MVU framework utilizes no video level training for any of its components and surpassed multiple approaches that rely on video-language training.
@@ F.3 Detailed Prompting Example
Appendix G Open-Ended Video Question Answering
Appendix H Longer Video Question Answering
While established long video benchmarks used as the key evaluations in numerous prior work (Wang et al., 2025; 2024c; Min et al., 2024; Park et al., 2024; Zhang et al., 2023a; Kahatapitiya et al., 2024; Wang et al., 2023b) limit to roughly 1-3 minute long videos, some newer datasets include even longer videos (Wu et al., 2024). We explore such even longer videos by evaluating our method on the LongVideoBench dataset (Wu et al., 2024).
We select Phi-3-Vision-Instruct (Abdin et al., 2024) as our baseline since it is the best performing model we can replicate within our compute budget. We note that larger sized models using significantly larger context lengths are difficult to replicate within academic compute restraints. Results using this baseline from Abdin et al. (2024) and our MVU framework integrated over it are presented in Table A.7. MVU gains clear performance gains in this longer video dataset.
Table A.7: LongVideoBench Evaluation: We integrate MVU with the baseline from Abdin et al. (2024) and highlight the additional performance improvements achieved by our MVU framework.
Method Acc (%)
Phi-3-Vision-Instruct (Abdin et al., 2024) 49.7
Phi-3-Vision-Instruct + MVU 50.4
@@ Appendix H Longer Video Question Answering
Appendix I Additional Ablations
Therein we investigate a new motion focused dataset, Something-Something-v2 (Goyal et al., 2017) (SSv2), only for this single ablation. The SSv2 dataset focuses on motion-based category discrimination, providing an ideal evaluation to measure the usefulness of our object motion trajectory modality. We benchmark on a subset of this dataset following (Lin et al., 2023b) and report these results in Table A.8. Our results while exceeding their performance also indicate the clear performance gains obtained when injecting the object motion trajectory modality into our MVU framework.
We also provide an ablation on frames used with our MVU framework in Table A.9. Increasing the number of frames leads to improved performance in contrast to some prior works (Mangalam et al., 2023) highlighting how our information fusion pipeline allows better utilization of the LLM context length. Additionally, the lightweight object detector and tracker used in MVU allows scaling the number of frames with a lesser increase in inference time.
Table A.9: Frame Count Ablation: We illustrate the importance of appropriate prompting when combining with likelihood selection, specifically for long video QnA tasks. Top-1 accuracy (%) is reported on EgoSchema subset (ES-S) and NextQA test set (NQA-T).
Method Frames EgoSchema-S Time (s)
MVU 16 60.3 2.42
MVU 32 60.4 2.48
MVU 64 60.4 2.60
MVU 128 61.2 2.81
Appendix J Tokenization in LLMs
Most modern LLMs utilize Byte-Pair Encoding (BPE) tokenization (Sennrich et al., 2015) to convert natural language into discrete tokens that are mapped to vocabulary embeddings. This process is learned from language itself and the resulting tokenization may sometimes break complete words into pieces (e.g. example $\rightarrow$ ex-am-ple). Given our utilization of logits extracted from an LLM forward pass, we note that each logit corresponds to a single token, which may at times be the embedding of some meaningless word piece. However, our calculation of a joint likelihood across a sequence of tokens ensures a meaningful process, which is validated by our strong results.
Appendix K LLM Context Length
Using LLMs for long video understanding has proven successful (Zhang et al., 2023a; Wang et al., 2024c) but handling long context lengths remains a key issue (Mangalam et al., 2023; Park et al., 2024), often leading to lower performance when additional frame information is provided to the LLM. This draws importance to frame selection, but we argue that alternate forms of information bottlenecks can also provide improvements, often complementary to frame selection.
In MVU, instead of naively collecting all information within a frame, we only collect object centric spatial and motion information, allowing to process more frames at a fixed context length. In other words, MVU information extraction from multiple frames can be viewed as an alternative to frame selection. This is because our object information extraction indirectly acts as an information bottleneck similar to frame selection. For frames without objects of interest, no information is extracted. For multiple frames containing the same object (identified by our object tracker), the repetitive information is removed. This resembles the idea of selecting useful information from multiple frames.
In fact, when comparing the average token length for a similarly performing baseline (implemented under identical settings using a common LLM), we use less tokens (context length) to achieve similar results. We show these results in Table A.10.
Table A.10: Context Length Comparison: We compare the context length used (i.e. number of tokens) to achieve similar results with LLoVi (Zhang et al., 2023a) as opposed to our MVU. We achieve better performance utilizing less tokens.
Method Average Tokens ES-F (%)
LLoVI 1940 33.5
MVU 1124 37.6
Appendix L Additional Baselines
We implement a multi-frame baseline directly using LLaVA-1.5 (Liu et al., 2023a) with no video level training. These results are reported in Table A.11. Results indicate that directly adding multiple frames to a VLM with no video level training does not lead to improved performance. Similar trends are observed in prior work (Kahatapitiya et al., 2024). These findings highlight the importance of careful per-frame information extraction and cross frame information fusion proposed in our MVU.
Table A.11: Multi-Frame LLaVA Baseline: We implemented multi-frame variants of LLAVA (Liu et al., 2023a) with no video level training. Results indicate that without any video level training such naive extension does not lead to results improvements.
Method Frames ES-S
LLaVA 1 55.8
LLaVA 8 53.4
LLaVA 16 46.2
LLaVA 32 40.2